OpenStudy (anonymous):

lim as theta approaches 0 (cos theta-1)/sin theta

OpenStudy (luigi0210):

Multiply by the conjugate?

OpenStudy (luigi0210):

And welcome to openstudy btw :)

OpenStudy (anonymous):

ohh ok so I would then multiply by cos theta + 1 ??

OpenStudy (anonymous):

& thanks :))

OpenStudy (luigi0210):

Yup!

OpenStudy (anonymous):

I was thinking multiplying by \(\dfrac{x}{x}\) would be simpler.

OpenStudy (luigi0210):

I doubt you know L'hopital? And won't they give you the same thing, Siths?

OpenStudy (anonymous):

No I don't no L'hopital ? & umm let me try

OpenStudy (anonymous):

Yes, but not if you separate the limits: \[\lim_{x\to0}\frac{\cos x-1}{\sin x}\cdot\frac{x}{x}=\color{blue}{\lim_{x\to0}\frac{\cos x-1}{x}}\cdot\color{red}{\lim_{x\to0}\frac{x}{\sin x}}\]

OpenStudy (luigi0210):

Sith, I don't think she's that far in calc :3 Let's just stick to the basics for now

OpenStudy (anonymous):

lol i'm in cal 1

OpenStudy (anonymous):

These limits are usually presented early on in the course. The blue is equivalent to 0, the red to 1.

OpenStudy (anonymous):

oh so the answer would be 0 right? because they are being multiplied together

OpenStudy (luigi0210):

Pretty much, right?

OpenStudy (anonymous):

Yes. Have you learned these limits yet?

OpenStudy (anonymous):

not really :/ my professor briefly went over limits using trig functions so i'm confused :(

OpenStudy (luigi0210):

Don't worry, you got us :3

OpenStudy (anonymous):

lol thanks :) I wasn't sure if this website was going to help :O

OpenStudy (luigi0210):

Trust me, it does. You just got to know the right people around here. @SithsAndGiggles is a great help around here

OpenStudy (jdoe0001):

*cough* khan academy has great videos *cough*

OpenStudy (luigi0210):

Khan Academy? Pfft >.>

OpenStudy (anonymous):

I have to go but I have some more questions where will I find you guys when I come back?

OpenStudy (luigi0210):

Probably on here still xD

OpenStudy (anonymous):

lol ok hopefully!

OpenStudy (luigi0210):

You can also call over users by using the "@" and putting their names.

OpenStudy (luigi0210):

Like @kari.lynee